### v1

@@ 8 Conclusion and remarks
This paper introduces SpoofCeleb, a dataset for SDD and SASV based on in-the-wild data. To create a dataset that incorporates real-world conditions, we developed a fully automated pipeline to process the VoxCeleb1 dataset, making it possible to use it for training TTS systems. The resulting bona fide data, TITW, also serves as a standalone dataset to encourage the training of TTS models using in-the-wild data. We further trained 23 TTS systems, partitioning TITW and the TTS systems into the SpoofCeleb dataset, which includes training, validation, and evaluation partitions. Protocols were defined to train and test both SDD and SASV models, and baseline systems for SDD and SASV were established, trained, and evaluated.
While numerous datasets exist for SDD, many are limited in scale or speaker diversity, which has hindered research on single SASV models. We hope SpoofCeleb serves as the first dataset with enough data to effectively train single SASV systems. Yet, SpoofCeleb has its limitations. In the experiments, some spoofing attacks are shown to be less challenging, as the wild nature of the TITW data complicates the training of robust TTS systems. Future work will focus on advancing TTS training techniques that can better leverage this challenging in-the-wild data.

### v2

@@ 8 Conclusion and remarks
This paper introduces SpoofCeleb, a dataset for SDD and SASV based on in-the-wild data. To create a dataset that incorporates real-world conditions, we used a fully automated pipeline to process the VoxCeleb1 dataset, making it possible to use it for training TTS systems. We further trained 23 TTS systems, partitioning TITW and the TTS systems into SpoofCeleb, which includes training, validation, and evaluation sets. Protocols were defined to train and test both SDD and SASV models, and baseline systems for SDD and SASV were established, trained, and evaluated.
While there are numerous SDD datasets, many are limited in scale or speaker diversity, which has hindered research on single SASV models. We hope SpoofCeleb will serve as the first dataset with enough data to effectively train single SASV systems. Yet, SpoofCeleb has its limitations. In the experiments, some spoofing attacks are shown to be less challenging, as the wild nature of the TITW data complicates the training of robust TTS systems. Future work will focus on advancing TTS training techniques that can better leverage this challenging in-the-wild data.